OpenStudy (anonymous):

Identify the correct slope and y-intercept of the equation 5x - 3y = -6 slope = -5; y-intercept = (0, -6) slope = five thirds; y-intercept = (0, 2) slope = -three fifths; y-intercept = (0, 2) slope = five thirds; y-intercept = (0, -6)

OpenStudy (anonymous):

i think its B

OpenStudy (anonymous):

slope of the equation is =5/3 for y intercept put x=0 y=2 (0,2) option b is more appropriate